OpenStudy (anonymous):

(2+1) to the power of 3/3 to the power of 2 im going to draw

OpenStudy (anonymous):



OpenStudy (anonymous):

3.

OpenStudy (anonymous):

\[\frac{(2+1)^3}{3^2}=\frac{3^3}{3^2}=3\]

OpenStudy (anonymous):

3